OpenStudy (anonymous):

2(x - 3) + 2(x - 7) = 40

OpenStudy (uri):

Solve the parentheses first 2(x - 3) + 2(x - 7) = 40 DO you know how to? @Challiexoxo

OpenStudy (cggurumanjunath):

2x-6+2x-14=40 4x-20=40 4x=40+20=60 x=60/4=15

OpenStudy (anonymous):

Thanks. And yes I do, I was just confused with this problem

OpenStudy (uri):

Do you understand? Do you want to learn how to do it?

OpenStudy (anonymous):

Yeah that'd be great!

OpenStudy (uri):

Hm so first solve the parentheses,What do you get? 2(x - 3) + 2(x - 7)

OpenStudy (anonymous):

2x-6 + 2x-14?

OpenStudy (uri):

Correct. Now, 2x-6 + 2x-14=40 Bring all numerials at one side,and bring all x values at one side.

OpenStudy (anonymous):

2x+2x= 40-6-14???

OpenStudy (uri):

2x-6 + 2x-14=40 2x+2x-6-14=40

OpenStudy (uri):

Now add all the x values and shift all numericals after *=* sign..remember when we shift sign changes.

OpenStudy (anonymous):

I'm confused.

OpenStudy (uri):

We can do one thing more: 2x+2x-6-14=40 4x-6-14=40 Add 6 and 14 both sides 4x-6-14=40 +6 +14 +6 +14 =========== 4x=60

OpenStudy (uri):

Now divide 60 with 4 what do you get?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

Ahh makes so much more sense. Thanks!!

OpenStudy (uri):

You're welcome :) @Challiexoxo